OpenStudy (anonymous):

-3y+6=-54 solve the problem step by step

OpenStudy (uri):

Bring 6 to the right side,sign changes. -3y+6=-54 -3y=-54+6

OpenStudy (anonymous):

-60/-3 = 20

OpenStudy (uri):

-54-6*

OpenStudy (saifoo.khan):

How does the sign change?

OpenStudy (uri):

Sign changes when you shift values to other side.

OpenStudy (anonymous):

this one was a easy one

OpenStudy (saifoo.khan):

Incorrect.

OpenStudy (uri):

-3y=-54-6 -3y=-60

OpenStudy (anonymous):

sign changes with add and subrtraction

OpenStudy (saifoo.khan):

Pakistani teachers skip one step and teach us wrong too.

OpenStudy (anonymous):

-3y+6=-54 SUBTRACT 6 FROM BOTH SIDES -3y+6=-54 -6 -6 --- ---- -3y + 0 = -60 -3y = -60 DIVIDE BOTH SIDES BY -3 -3y = -60 ---- ---- -3 -3 y = 20

OpenStudy (saifoo.khan):

-3y+6=-54 We actually add '-6' on both sides. Making it: -3y+6-6 = -54 - 6 Then in the next step: -3y = -60

OpenStudy (saifoo.khan):

There^ The apple guy did it.

OpenStudy (uri):

Well both ways are correct..

OpenStudy (saifoo.khan):

Incorrect, again.

OpenStudy (uri):

I did the way i'm taught..and i was right.

OpenStudy (uri):

-3y+6=-54 -3y=-54-6 -3y=-60 y=-60/-3 y=20.

OpenStudy (saifoo.khan):

We, here, aren't taught that step.

OpenStudy (saifoo.khan):

You actually change the sign when you add the same term with different signs on both sides.

OpenStudy (anonymous):

@uri what do you mean by "sign changes"?? I don't seem to understand that concept.

OpenStudy (uri):

When we shift a value before or after = sign,the sign changes.That's what i'm taught like this, 4x-25=23-2x 2x+4x=23+25

OpenStudy (saifoo.khan):

Uri, that's what i'm saying. We aren't taught that step which i showed you. I too didn't know about it before.

OpenStudy (uri):

Okay now i know...

OpenStudy (uri):

But this way also we get the answer correct...no?

OpenStudy (saifoo.khan):

The point is not to get the correct answer. The point is to know HOW to arrive to that answer.